OpenStudy (anonymous):

2x^2-3x-15=5 2x^2-3x-15=5 @Mathematics

jimthompson5910 (jim_thompson5910):

2x^2-3x-15=5 2x^2-3x-15-5=0 2x^2-3x-20=0 2x^2-8x+5x-20=0 2x(x-4)+5(x-4)=0 (2x+5)(x-4)=0 2x+5=0 or x-4=0 2x=-5 or x=4 x=-5/2 or x=4

OpenStudy (anonymous):

Are you the guy that helped me last time?! :D

jimthompson5910 (jim_thompson5910):

with history yeah

jimthompson5910 (jim_thompson5910):

how did that go

OpenStudy (anonymous):

Very good, thanks! I thought you helped me with math too.. hmm.

jimthompson5910 (jim_thompson5910):

maybe...not sure, but you're probably right

OpenStudy (anonymous):

Are you a teacher or student or..?

jimthompson5910 (jim_thompson5910):

student learning to become a teacher so both in a way

OpenStudy (anonymous):

Ah, so in college?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

x^2+2x-1=2 x^2+2x-3=0 Not sure where to go from there

jimthompson5910 (jim_thompson5910):

find two numbers that BOTH a) multiply to -3 AND b) add to 2 These two numbers are 3 and -1 So we break up 2x into 3x-1x to get x^2+2x-3=0 x^2+3x-1x-3=0 x(x+3)-1(x+3)=0 (x-1)(x+3)=0 I'll let you take over from here